Some systems are experiencing issues

About This Site

Maintenances and incidents regarding our services, networks and datacenters can be followed here.

Scheduled Maintenance
[DC3] Maintenance of fire suppression system

[DC3] Maintenance of fire suppression system

Maintenances:

-Control VESDA system

-Control manual triggers

-Control of sound and visual alarms

-Control Actuated Devices of Safety

-Control alarm reports

-Controle valves fire suppression (outside datarooms)

-Control smoke extraction and extraction engine.

Impact : No impact for customers

Start : 9-Dec-2019 0500Z 0600LT

Duration : 14 Days

[DC3] - Curative maintenance on battery of UPS

[Curative maintenance on battery of UPS]

We will do the curative maintenance of the inverter of electrical chain A. Operations will be performed by our maintainer under our permanent supervision. The downtime of an UPS is estimated to take 30 minutes.

The operation will be done by UPS and will have no impact on your equipments housed in our Datacenter.

Location : DC3 DATACENTER

Impact : None

Start : 11-Dec-2019 0730Z 0930LT
End : 11-Dec-2019 1530Z 1730LT

Duration : 1 day

Past Incidents

Sunday 15th September 2019

RPN AMS1/DC3, scheduled 2 months ago

Preventive maintenance over the RPN network between AMS1 and DC3

Network supplier must perform maintenance on his Network.

Maintenance window is from September 14, 2019 at 11:00 pm UTC to September 15th, 2019 at 5:00 am UTC.

The expected Impact length is 6 hours at the maximum, during the maintenance window

Impact : Possible downtime on the RPN network between AMS1 and DC3

Wednesday 11th September 2019

DC2 SCW - PF10 - Optical maintenance, scheduled 2 months ago

[SCW DC2 -PF10] Optical replacement + fiber cleaning

Operation completed successfully - Closing the status

Network glitch during 2 minutes.

Start : 11-Sept-2019 0730Z 0930LT
Start : 11-Sept-2019 1030Z 1230LT Operation completed successfully

Duration : 2 minutes

Tuesday 10th September 2019

RPN DC5-A Room 120 C13 & C15 Optical Maintenance, scheduled 2 months ago

[RPN DC5-A Room 120 C13 & C15] Optical replacement + fiber cleaning

Operation completed successfully - Closing the status

RPN switch down during 2 minutes.

Start : 10-Sept-2019 0900Z 1100LT

Start : 11-Sept-2019 1030Z 1230LT Operation completed successfully - Closing the status

Duration : 2 minutes

DC3: Room 32 Rack 10 & room 43 rack16 Optical Maintenance, scheduled 2 months ago

[DC3: Room 32 Rack 10 & room 43 rack16 ] Optical replacement + fiber cleaning

Network glitch during 2 minutes.

Start : 10-Sept-2019 0730Z 0930LT

Duration : 2 minutes

Thursday 5th September 2019

AMS1 [AMS] Room Hall 2 E-E Rack E14 - Power delivery issues

A rack in AMS1 is having power delivery issues, all servers in the rack are unreachable.

===================

09.05.19 1620Z (1820LT)

Issue has been escalated to local team

09.05.19 1815 (2015LT)

The product team has identified the root cause and working on a fix

09.05.19 1830 (2030LT)

Incident has been solved and servers are coming back up progressively.

Mail [HOSTING] Mail Connexion Issue

Our teams identified an issue regarding webmail connection. We are currently investigating the issue and will update this status as soon as we have more details to share.

===================

05/09/19 1230 (1030LT)

The product team has identified the root cause and working on a fix

05/09/19 1730 (1530LT)

Issue has been fixed now, if you are still experiencing connection issues please contact our support team via ticket.

Wednesday 4th September 2019

DC3 - Curative maintenance, scheduled 2 months ago

[DC3] Replacement of source switching in pumps cabinet

GF1substation A

No impact on users

Start : 4-Sept-19 0730Z 0930LT

Duration : 8 hours

Web Website and FTP unavailability

We have detected issues with several website on web hosting platforms. Our teams are currently working on solving the problem.

Impact: Some website / FTP may be unavailable.

===================

09.04.19 - 1440Z (1640LT)

Issue has been fixed and websites are now working normally.

09.04.19 - 1300Z (1500LT)

FTP Access has been restored.

09.04.19 - 1202Z (1402LT)

The issue is currently being investigated.

09.04.19 - 1200Z (1400LT)

Issue has been escalated to local team

Object Storage Error 403 on requests

We have problems with object storage requests, a 403 error may appear during your requests. Our teams are currently working on solving the problem.

Impact: list your buckets may be impossible, accessing your objects returns 403 errors.

===================

09.040.19 - 1330Z (1530LT)

Issue has been fixed and buckets are now working normally

09.04.19 - 0800Z (1000LT)

The issue is currently being investigated.

09.04.19 - 0600Z (0800LT)

Issue has been escalated to local team

Tuesday 3rd September 2019

No incidents reported

Monday 2nd September 2019

No incidents reported

Sunday 1st September 2019

No incidents reported

Saturday 31st August 2019

No incidents reported

Friday 30th August 2019

No incidents reported